2017/18 Season HUB Weekly Farm Update as at 9th January 2018
Hub fodder beet just on canopy closure 9/1/18
Farm manager employed for the 2018/19 season!
We were fortunate to have some very well qualified applicants for the farm manager role for next season and we are thrilled to announce that Shane Griffin has accepted the role starting June 1st.
Shane is currently managing a 500-cow property in Northern Southland and is looking forward to the new challenge The Hub will provide! We will involve Shane in the management decisions that need to be made as we set up the 4 farmlets for next seasons research. Welcome on board Shane!
This is a good time to acknowledge our present farm manager Jason. He has made an invaluable contribution to The Hub this season with his experience in managing research farm activities in the Waikato. Jason was on loan to us from DairyNZ for the season and has elected to return to the Waikato in June.
Conditions for the Wallacetown area remain dry with an impact on grass and crop growth and production. We have a BCS assessment planned to determine the impact of the feeding conditions on body condition score.
All the crops have had a weed spray, but the kale crops require another this week as the dry conditions restrict crop growth. With a little rain last week, we were able to get Urea on the crops with more booked to go on in February if conditions permit.
Critical Targets this week
- Kale weed spray
- Weigh R1 heifers
- Bull out of the R2 heifers and the cows at the end of this week (11 weeks)
- Farmwise consultant visit
- Prepare and revise for our next Farmtune meeting next week
Herd & Production
Four herds split evenly on age, BW / PW and on the winter crop trial treatment groups to ensure the herds are as even as possible. Each herd allocated a farmlet corresponding to their herd tag colour Green, Blue, Yellow and Pink. Farmlets have paddocks allocated so each herd has equal walking distance from the shed and the same proportion of each soil type and equal proportions of pastures in the FVI trial (forage value trial – refer web site section on research).
Staph Aureus has been identified in approximately 40 cows at The Hub. A high bulk SCC and cows not reacting quickly to treatment encouraged investigation leading to this diagnosis. Infected cows have drafted out and are being milked last to prevent the spread of this infection.
Vet South has completed a ‘Milking Assessment’ and have provided a mastitis treatment and management plan for the team to follow. RMT tests confirm which cows have clinical mastitis and these cows are kept out of the vat whilst clean cows are milked into the vat. This has lifted bulk SCC from 98,000 to 120,000 approximately.
The treatment plan so far is –
Milking Assessment results
- Machine check completed, no vacuum issues
- Some cups were twisted when cows were being milked – could have resulted in un-even milk out. Team ensuring rubberwear kept hanging correctly at milking.
- Teat spray appears to be getting good coverage
- Test done on teat spray concentration, all good
- Hygiene – some cows had dirty tails these have been trimmed to prevent poor hygiene
- More leg bands for all 3-titters to avoid cupping blind quarters
- Check for stray voltage to be completed as some cows appeared unsettled when the cups were removed
- Plan a culling guide and dry cow strategy with Sally at Vet South to eliminate the problem at the end of the season.
Production has dropped from 1.6kgMS/cow/day to 1.55kgMS/cow/day as pasture supply has declined with the dry conditions affecting growth. The number of cows on OAD to manage BCS within the herds is stable but we have a new BCS assessment due this week so this number on OAD may rise.
11 cows in the Staph mob have been identified to cull as they have either –
- Been confirmed with Staph Aureus
- Are 8 years of age or older therefore less than a 30% chance of cure
- Have poor udder conformation
For now, the remainder will be milked until the end of the season, into the vat when clear of clinical mastitis then dried off in conjunction with vet advice from Sally at VetSouth.
Table 1: Herd & Production
|Herds on Milking Platform||Unit||Value|
|Milkers - TAD||cows||585|
|Milkers - OAD||cows||59|
|Sick mob (Staph, lame etc – approx. half in vat)||cows||42|
|Culls / deaths||cows|
|Current Stocking Rate (platform)||cows/ha||2.7|
|Production (to the factory) (21/11/17)|
|Total Milksolids (YTD)||kgMS||149,867|
|Total Milksolids (MTD)||kgMS||7,105|
|Total Milksolids (last 7 days)||kgMS||7,105|
|Average per cow daily production (last 7 days)||kgMS/cow/day||2,035|
|Per ha production (last 7 days)||kgMS/ha||4.02|
|Other (to the factory)|
|Protein:Fat Ratio (7-day avg.)||%||0.81|
|Milk Urea (7-day avg.)||mg/dl||22.61|
|SCC (7-day avg.)||cells/ml||102,000|
|Fat Evaluation Index (FEI)||A|
|Fonterra 17/18 Forecast - as at 27 July 17||$/kgMS||$6.40|
The four herds are to remain on their allocated farmlets at all times. The 2017/18 season aim is to feed all cows in a similar manner all season. Any differences to be recorded. OAD (BCS of 4 or less) cows can be fed extra supplement in the shed as required by the farm team but will remain with their allocated herds for research purposes. Concentrate in shed is a 50:50 mix of PKE and barley.
Average pasture cover continues to drop with residuals holding at the existing base where the topping mower cut to at approx. 1500 – 1600kgDM/ha.
All cows have been receiving 1.5kg/c/day PKE/Barley 50:50 mix this last week to hold residuals, round speed and cow BCS.
We have a stack of pasture silage remaining on farm but at lower quality (9MJME) we are looking to use this in the autumn when a poorer quality feed will not have as much impact. The length and severity of the current dry conditions may alter this thinking.
Table 2: Feed Offered
|Milkers – Pasture offered||kgDM/cow/day||16|
|Milkers - Fodder Beet||kgDM/cow/day||0|
|Milkers - Pasture Silage||kgDM/cow/day||0.00|
|Milkers - Concentrates - TAD Cows||kgDM/cow/day||1.5|
|Milkers - Concentrates - OAD Cows||kgDM/cow/day||1.5|
Land, Pasture, Crop & Silage Management
Total area is 327ha(eff), 32.5ha(eff) young stock block and 294ha(eff) platform. Winter crop is grown on both blocks for all stock. Crop rotation is two years in crop then re-grass. PGG Wrightson grass cultivars used are Rely, Platform and Excess, all diploids. Where target residuals are missed, topping after grazing is the tool the farm team use as needed.
All the R1 heifers remain on the platform until the stock yards are built on the young stock block. From the 12/1/18 all R2 heifers will be on the young stock block.
Table 3: Land, Pasture, Crop & Silage Metrics
|Pre - Graze cover target||kgDM/ha||2700|
|Post - Graze cover target||kgDM/ha||1600|
|Average Pasture Cover (avg. over 4)||kgDM/ha||2120|
|Growth Rate (avg. over 4 farmlets)||kgDM/ha/day||39|
|Rotation Length (avg. over 4 farmlets)||days||28|
|Area Available for grazing (milkers)||ha||253|
|Regrassing Area (out of rotation)||ha|
|Baleage/Silage Area (out of rotation)||ha||0|
|Crop Area (out of rotation)||ha||46.4|
|Total platform effective||ha||300|
|Young Stock (not included in rotation)||ha||29.0|
Table 4: Fertiliser & Nitrogen Metrics
|Nitrogen (Urea only)(last 7 days)||Unit|
|Area – Crop area 46ha||ha||46|
|Rate – Urea @100kg/ha||kg/N/ha||100|
|Fertiliser (last 7 days)||Unit|
|Rainfall (total last 7 days)||mm||5.6|
|Soil Temp (7-day average) @10am||°C||18|
- 90% submission rate 3 weeks) (91% achieved)
- 78% 6-week in calf rate (Sitting on 68% after 3 weeks)
Bulls will be taken out of both R2’s and cows by the 12th January (21st October last day calving)
The farm team report little bulling action in either cows or heifers, so the decision has been made to pull the bull out a week early. This also coincides with a new team member starting next week and the wish to ensure she starts safe with us. We leave the bulls behind in the paddock rather than have them on the lanes and with 4 herds to manage this with twice a day our ‘cow chaser’ job is not straight forward!
Table 5: Reproduction Metrics
|Planned start of Mating (PSM)||Unit|
|Mixed Aged Cows||Date||29-Oct-17|
|2016 Born Heifers||Date||21-Oct-17|
Young Stock Management
The other half R2 heifers are due to be weighed this week and drenched
Research Update (fortnightly)
No update this week. The DairyNZ research tech team will complete a herd test and a herd BCS this week.
The SDH welcome research proposals for any sampling or research on the SDH, these are assessed by the Research Advisory Committee (RAC). Just send your request or ask for information via firstname.lastname@example.org
For more information check out the DairyNZ link:
- 8:2 & 8:3 roster
- The last of the farm team are completing a few days off now. The challenge will be to find some more time to take some annual leave over the summer and autumn months and enjoy the weather while we can
- As already mentioned we welcome the announcement of a new farm manager for the 2018/19 season Shane Griffin to start on June the 1st
- We look forward to Ashlee Kidd joining as our newest team member next week
- Advertising continues for another permanent employee to lift our team to 5 permanent farm staff for the rest of this season and into the next. Interviews starting soon.